OpenStudy (anonymous):

Medal Find the area of a square tile that is 3 feet long on each side. A. 3 ft² B. 6 ft² C. 9 ft² D. 12 ft²

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello

OpenStudy (studygurl14):

Area of Square = (length of side)^2

OpenStudy (misty1212):

there must be something wrong in the question the area of a square is the square of the side that is why it is called "squaring' there is no \(\pi\) involved at all

OpenStudy (studygurl14):

Why must something be wrong ? Looks fine to me.

OpenStudy (misty1212):

oh doe that is ft i thought it was \(\pi\) nvm

OpenStudy (anonymous):

I dont see a pi involved here.

OpenStudy (misty1212):

just do \(3\times 3\) and you will have it

OpenStudy (anonymous):

Ah thank you! Im a bit rusty on my math ^_^;

OpenStudy (studygurl14):

:)